Find the equation passing through (1, -3) and perpendicular to 4y + 2x = 1
Answer by ikleyn(52776)

.
Find the equation passing through (1, -3) and perpendicular to 4y + 2x = 1
~~~~~~~~~~~~~~~~~~~~~~~

Perpendicular to  4y + 2x = 1  is  2y - 4x = c,  where "c" is some constant value.

(We swap x and y in the original equation and change the sign one time).


To provide  " through (1,-3) ",  we should select/find  "c"  in appropriate way.


For it, we substitute the coordinates x=1, y= -3  into equation 2y - 4x = c.
You get then

    2*(-3) - 4*1 = c,

or

    -6 - 4 = c.

Thus c = -10.


So, the final equation is

      2y - 4x = -10.     ANSWER

Solved.

------------------

For this subject, read and learn from the lesson
    - Equation for a straight line perpendicular to a given line and passing through a given point
in this site and find there the complete explanation on how to do it.

At this site,  there is a group of lessons related to this class of problems
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point (*)
    - Advanced problems on finding equations for straight lines

The most relevant to your current problem is the lesson marked  (*)  in the list.
So start from this lesson.

But if you want to learn the subject in all its aspects,  you need to learn all these lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.